OpenStudy (anonymous):

One savings account begins with $1000 and earns 6% interest, compounded monthly. Another account begins at the same time with $1200 and earns 4% interest, compounded yearly. How long will it take for the two accounts to reach the same value? Give an answer in years, rounded to the nearest whole number.

OpenStudy (anonymous):

How do I determine when the two accounts reach the same value?

OpenStudy (anonymous):

A1 = 1000(1+0.06/12)^12t A2 = 1200(1+0.04)^t set A1=A2, take logarithms and solve for t, the time in years

OpenStudy (anonymous):

before you take logs of A1 and A2, do you first divide? @douglaswinslowcooper

OpenStudy (anonymous):

I am stuck at the step where you take logs of the two accounts

OpenStudy (anonymous):

Either way should work. Dividing by 1000 simplifies a bit. (1.005)^12t = 1.2(1.04)^t 12 t log (1.005) = log(1.2) + t log(1.04) t [12 log(1.005)-log(1.04)] = log 1.2 t[0.026 -0.017] = 0.079 t = 8.8 years = 9 years check 1000(1.005)^108 = 1714 1200(1.04)^9 = 1708 close enough for our purposes

OpenStudy (anonymous):

Glad to have been of help.

OpenStudy (anonymous):

@douglaswinslowcooper Can you explain the first step when you simplify?

OpenStudy (anonymous):

(1.005)^12t = 1.2(1.04)^t I see where you log both sides

OpenStudy (anonymous):

about tree fiddy

OpenStudy (anonymous):

1000/1000=1 and 1200/1000=1.2 if that is the step you mean. I divided both sides by 1000.

OpenStudy (anonymous):

the second one you would want to divide by 1200... right?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

dividing both sides of the equation by 1000 just to make the numbers shorter